OpenStudy (sofiaduque):

simplify the expression -(3x-4)

OpenStudy (skullpatrol):

property of the opposite of a sum $$-(a+b)=(-a)+(-b)$$ the opposite of a sum real numbers is equal to the sum of the opposites of the numbers

OpenStudy (skullpatrol):

When you find the opposite of a sum or a difference, you change the sign of each term of the sum or difference.

OpenStudy (skullpatrol):

Also since -a = -1a -(3x-4) = -1(3x-4)

OpenStudy (kevin):

you need to multipy '-' to the number in parenthesis.